OpenStudy (technodynamic):

Teach me how to solve piecewise functions please step by step: f(x) ={x^2+2x if x<=-1 x if -1 x< x <=1 -1 if x >1 Use f(-4), f(-3/2)

OpenStudy (anonymous):

when x=-4, x is less than -1

OpenStudy (anonymous):

so you sub into the formula f(x) = x^2+2x

OpenStudy (anonymous):

f(-4) = (-4)^2 +2(-4) = 8

OpenStudy (anonymous):

when x=-3/2 , x is still less than -1, so you use the same expression

OpenStudy (anonymous):

f(-3/2) = (-3/2)^2 + 2 (-3/2) = 9/4 -3 = 9/4 -12/4 = -3/4

OpenStudy (technodynamic):

so when I use the input, i would first look to the right to see which is less than or equal to?